-32y to the 2nd -24y=0
i need help on solving this problem
6h to the 2nd =3h
-42z to the 2nd =14z
-12p to the second= -8m

Answer by stanbon(75887)

-32y to the 2nd -24y=0
---
-32y^2 - 24y = 0
-8y(4y+6) = 0
y = 0 or y = 3/2
-----------------------------
i need help on solving this problem
6h to the 2nd =3h
---
6h^2 = 3h
6h^2-3h = 0
3h(h-1) = 0
h = 0 or h = 1
------------------------
-42z to the 2nd =14z
---
-42z^2 - 14z = 0
-14z(3z+1) = 0
z = 0 or x = -1/3
-----------------------
-12p to the second= -8m
-12p^2 + 8m = 0
-4p(3p-2) = 0
p = 0 or p = 2/3
=========================
Cheers,
Stan H.